x-3y=2
3x-4y=-4
Cant solve this using substitution method.
any help would be appreciated
thanks

Answer by jim_thompson5910(28595)

Step 1: Solve x-3y=2 for x to get x = 3y+2


Step 2: Move onto 3x-4y=-4 and replace 'x' with '3y+2' (this is possible since x = 3y+2, ie, they are the same thing). So this substitution gives us 3(3y+2)-4y=-4


Step 3: Solve for y in the equation 3(3y+2)-4y=-4 (I'll let you do that)


Step 4: Use the answer from step 3 to find the value of x. You can use any equation, but x = 3y+2 is the easiest